OpenStudy (anonymous):

13+4a=9 please show steps also

OpenStudy (anonymous):

Hi, you need to do the same procedure as in the last exercise.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

substract 13 from both sides of the equation, and then divide both sides of the equation by 4.

OpenStudy (anonymous):

but.. idk really how..

OpenStudy (anonymous):

look: \[13+4a=9\]

OpenStudy (anonymous):

\[13-13+4a = 9-13\]

OpenStudy (anonymous):

\[4a=-4\]

OpenStudy (anonymous):

\[\frac{4a}{4}=\frac{-4}{4}\]

OpenStudy (anonymous):

a = -1

OpenStudy (anonymous):

im starting to see this now

OpenStudy (anonymous):

i did.